A coin lying on a meterstick starts to slide when one end of the stick is raised 36.4 cm above the horizontal. (a) What is the limiting angle for friction? (b) What is the coefficient of starting friction?

1
Expert's answer
2022-04-26T14:56:32-0400

a)

"\theta=\arctan\frac lm=20°,"

b)

"\mu=\frac lm=0.364."